OpenStudy (happpily):

What is the equation, in standard form, of a parabola that contains the following points? (–2, –20), (0, –4), (4, –20)

OpenStudy (danjs):

You are given 3 points

OpenStudy (danjs):

The standard form for a parabola y = ax^2 + bx + c

OpenStudy (danjs):

you need to find a,b, and c

OpenStudy (danjs):

We can use those 3 points, to generate 3 equations with 3 unknowns, a,b,c and solve for those 3 unknowns

OpenStudy (danjs):

Using the first point. (-2, -20) -20 = a(-2)^2 + b(-2) + c

OpenStudy (danjs):

so we end up with 3 equations like this -20 = a(-2)^2 + b(-2) + c -4 = a(0)^2 + b(0) + c -20 = a(4)^2 + b(4) + c

OpenStudy (danjs):

-20 = 4a - 2b + c -4 = c -20 = 16a + 4b + c

OpenStudy (danjs):

so right off the bat you find c = -4

OpenStudy (danjs):

so you are left with -20 = 4a - 2b - 4 -20 = 16a + 4b - 4

OpenStudy (danjs):

Understand so far?

OpenStudy (happpily):

yeah i do :)

OpenStudy (danjs):

so solving those 2 equations with , a, and ,b...

OpenStudy (danjs):

You will find...a=-2 and b=4

OpenStudy (danjs):

so the parabola containing those 3 given points, in standard form is: y = ax^2 + bx + c

OpenStudy (danjs):

y = -2x^2 +4x - 4

OpenStudy (danjs):

if you want, you can put in the points and see if the equation holds true , as a extra check

OpenStudy (danjs):

for example the point (-2, -20) -20 = -2(-2)^2 + 4(-2) - 4 ???

OpenStudy (danjs):

-20 = -8 - 8 - 4 = -20 TRUE

OpenStudy (happpily):

ohhh okay i totally get it now thank you!

OpenStudy (danjs):

no prob, anytime